OpenStudy (anonymous):

s(t)=t^2 -2t+4 is the distance function. At which point does the direction of motion change?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

t=1 sec

OpenStudy (anonymous):

from your eyeballs. vertex of this thing is \[-\frac{b}{2a}=\frac{-2}{2}=1\] so don't let anyone tell you to take the derivative etc

OpenStudy (anonymous):

actually it is \[-\frac{b}{2a}=-\frac{-2}{2}=1\]